Translate the following statement into an algebraic equation. Let x represent the number. 5 more than 13 times a number is 8 times that same number.
Answer by funmath(2925)

Translate the following statement into an algebraic equation. Let x represent the number. 5 more than 13 times a number is 8 times that same number.
"5 more than" is 5+
"13 times a number" is 13*x=13x
"is" means =
"8 times that same number" is 8*x=8x
The equation is:
13x+5=8x
Happy Calculating!!!